# Symmetry and Topology in the Non-Hermitian Kitaev chain

### v1

@@ I Introduction
Non-Hermitian quantum mechanics has engaged physicists for many years from various research areas, e.g. mathematical physics PhysRevLett.80.5243; Bender_2007; 10.1063/1.532860; PhysRevLett.89.270401; Mostafazadeh_2003, complex space quantum mechanics BENDER1993442; PhysRevD.73.085002, and scattering theory MOISEYEV1998212. More recently, researchers have been interested in understanding the quantum dynamics of open quantum systems, i.e. systems coupled with an environment PhysRevLett.108.173901; PhysRevA.95.022117; RevModPhys.89.015001; PhysRevA.106.L011501; Rotter_2009; PhysRevLett.124.196401 for their potential application in quantum optics 10.1088/978-0-7503-6014-2; Carmichael1993 and quantum information theory RevModPhys.88.021002; LewisSwan2019. Moreover, non-Hermitian generalizations of lattice models provide an excellent testbed to study phenomena ubiquitous to non-Hermitian physics, such as complex spectra, the appearance of exceptional points and the non-Hermitian skin effect (NHSE) i.e., exponential localization of the eigenstates towards one edge (or boundary) PhysRevLett.121.086803.
@@
The Hatano-Nelson model PhysRevLett.77.570, a generalization of tight-binding Hamiltonian with non-reciprocal hopping amplitude is an example of a system that exhibits NHSE. An unusual feature of such systems is the high sensitivity of the spectrum to the boundary conditions. This manifests itself in the fact that under periodic boundary conditions, the spectrum is complex and the system is unstable, but surprisingly in a finite open chain, the spectrum is purely real. The eigenstates, instead of being Bloch waves, localize exponentially toward one edge, giving rise to the NHSE.
@@
In a set of seminal work PhysRevLett.121.086803; PhysRevLett.123.246801, it was shown that NHSE has a deep connection with topological properties of the system. It was shown that with non-reciprocal hoppings in the Su-Schrieffer-Heeger (SSH) model, the topological phase transition point in a finite open chain does not coincide with that of an infinite (or periodic) chain. The true transition point is governed by non-Bloch bulk-boundary correspondence, which is central to topological non-Hermitian systems. This discovery has inspired a slew of research at the intersection of non-Hermitian physics and topology PhysRevLett.123.170401; PhysRevLett.124.196401; Zhang_2022; PhysRevLett.125.126402; PhysRevLett.124.086801; Okuma_2023; RevModPhys.93.015005; PhysRevLett.121.026808; PhysRevLett.123.066404; PhysRevLett.129.070401; Xiao2020; PhysRevLett.124.056802; PhysRevX.13.021007; PhysRevLett.123.170401; PhysRevB.99.201103; PhysRevLett.125.226402; PhysRevLett.122.237601; PhysRevLett.125.180403; PhysRevLett.130.017201; PhysRevB.100.054301; PhysRevLett.124.250402; PhysRevB.109.024306; PhysRevLett.126.230402; Ghatak_2019; PhysRevLett.132.113802; Kawabata_2022; PhysRevB.101.195147; PhysRevX.8.041031; PhysRevB.55.1142; PhysRevX.9.041015.
@@
This naturally raises the question of whether such behavior is universal in one-dimensional systems. By considering a non-Hermitian Kitaev chain Kitaev_2001; Alicea_2012; PhysRevA.94.022119; Li_2023; Yan_2023; PhysRevResearch.5.L022046; https://doi.org/10.1002/pssb.202200448; PhysRevB.111.205432; Ardonne_2025 with non-reciprocal nearest neighbor hopping amplitudes ($t_{1}$ and $t_{2}$) and asymmetric superconducting pairing terms ($\Delta_{1}e^{-i\varphi_{1}}$ and $\Delta_{2}e^{i\varphi_{2}}$), see Fig. 1, we show that this is not generally the case. For a non-Hermitian Kitaev chain, cf. Fig. 1, the topological phase transition points for finite open chain coincides with that of an infinite (or periodic) chain. By performing a symmetry analysis, we show that particle-hole symmetry (PHS) always exists across the entire parameter regime and thereby protects the Majorana modes from NHSE. However, we show that particle and hole bulk modes exhibit an equal and opposite skin effect, which is also a consequence of PHS. Finally, we present the phase diagram for the non-Hermitian Kitaev chain in the complex chemical potential ($\mu$) plane.
@@ II Kitaev chain
The one-dimensional (1D) Kitaev chain exhibits topological phase transition with two distinct phases in the parameter space. It is a tight-binding chain with $p$-wave superconducting pairing terms and hosts Majorana edge states in the topological phase Kitaev_2001. We begin with a brief review of the 1D (Hermitian) Kitaev chain Alicea_2012 before investigating its non-Hermitian extension. The Hamiltonian with $L$ lattice points, written in terms of particle creation ($c_{j}^{\dagger}$) and annihilation ($c_{j}$) operators on the $j^{\text{th}}$ site is given as follows,
@@
where $t$ is the hopping amplitude, $\Delta$ is the superconducting gap, $\varphi$ is the superconducting phase and $\mu$ is the chemical potential. This system exhibits a topological phase transition 10.1093/acprof:oso/9780199227259.001.0001 as we tune the parameters ($t$ or $\mu$) of the Hamiltonian. The topological phase is marked by the existence of Majorana zero edge modes in a finite chain Kitaev_2001; Alicea_2012; Aghaee2025.
@@ III.1 Finite non-Hermitian Kitaev chain
where we have again assumed for simplicity $\mu,t_{1},t_{2}\in\mathbb{R}$ and $\varphi_{1}=\varphi_{2}=0$. Diagonalizing the Hamiltonian in Eq. (6) reveals a complex band structure, Fig. 2b,c and the presence of zero-energy modes in certain parameter regime, Fig. 2a.
@@
The band structure shows that the zero mode solutions only appear for certain values of the chemical potential $\mu$, see Fig. 2a, to be precise, the zero modes exist whenever we have $|\mu|<\frac{1}{2}(t_{1}+t_{2})$. This implies that the topological transition point derived using the spectrum of infinite chain, Eq. (III) coincides with the transition point obtained from the band structure of finite chain with OBC, cf. Fig. 2a-b. In order to gain further insight, we look at the wave functions for the finite chain with OBC. With non-reciprocal hoppings ($t_{1}\neq t_{2}$), one would naively expect to see the non-Hermitian skin effect (NHSE) PhysRevLett.121.086803. However, we find that the zero modes do not show NHSE and remain robust to non-reciprocity, Fig. 3a-c. The NHSE only appears when we consider wavefunctions corresponding to excited states, Fig. 3d. In the following subsections, Subsec. III.2 and Subsec. III.4, we present an explanation for the robustness of the zero modes to NHSE, Fig. 3a-c, the presence of skin effect in excited states, Fig. 3d.
@@ III.2 Generalized Brillouin zone
The NHSE is an exponential localization of wave functions and is the result of a modified Brillouin zone (BZ) in non-Hermitian systems PhysRevLett.121.086803. Simplest way to obtain the modified Brillouin is to make the substitution $e^{ika}\to\beta$ (note that $a$ is the lattice constant and is set to one), where $|\beta|$ is determined by the non-reciprocities (or the non-Hermiticity) in the hopping parameters PhysRevLett.121.086803; PhysRevLett.123.246801. Hamiltonian $H^{\text{NH}}$, Eq. (4) with the substitution $e^{ik}\to\beta$ becomes,
@@
From Eq. (9) we find that if we have $\beta_{0}$ as a solution, then $\beta_{0}^{-1}$ is also a solution. Hence, in general, we have four solutions to Eq. (9): $\beta_{1},\,\beta_{1}^{-1},\,\beta_{2},\,\beta_{2}^{-1}$. Additionally, one can show that the matrix, Eq. (8) can be null only at the boundary of the topological and trivial phase, hence the rank of that matrix is 1 at all points in the topological phase. Hence, corresponding to each $\beta$, there is only one zero mode eigenvector in the topological phase. Thus, the four roots of Eq. (9) will have four zero mode eigenvectors in the topological phase. Specifically, at $\mu=0$, we have $\beta_{2}=-\beta_{1}$. On physical grounds, we expect that the zero mode eigenvectors corresponding to $\beta_{1}(\beta_{1}^{-1})$ will be the same as $\beta_{1}(-\beta_{1}^{-1})$. Consequently, for $E=0$, the two solutions have inverse localization factors $\beta_{1}$ and $\beta_{1}^{-1}$, which means that the solution $\beta_{1}$ ($\beta_{1}^{-1}$) localizes to the right (left) Li_2023; Yan_2023; https://doi.org/10.1002/pssb.202200448. The total probability density of the zero-energy states is given by $|\psi_{0}^{(1)}|^{2}+|\psi_{0}^{(2)}|^{2}$, where $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ are the normalized zero modes corresponding to factors $\beta_{1}$ and $\beta_{1}^{-1}$ respectively. This reciprocal pairing eliminates the skin effect for zero-energy states.
The existence of the reciprocal solution pair can also be understood by the symmetries of the Kitaev chain. Particle-hole symmetry enforces solutions to be reciprocal pairs Li_2023. In subsec. III.3, we demonstrate that the particle-hole symmetry exists for arbitrary complex parameters in the Hamiltonian, thus protecting the Majorana edge modes, whenever they exist, for those parameter values. However, for any excited state (bulk mode), the solutions to $\text{det}(H^{\text{NH}}(\beta)-E)=0$ are not in a reciprocal pair. In fact, if $\beta_{1}$ is a solution corresponding to energy $E$, then $\beta_{1}^{-1}$ is a solution corresponding to energy E-E. Thus we expect an equal and opposite non-Hermitian skin effect for the particle and hole wavefunctions, cf. Fig. 4b.
@@ III.3 Symmetry analysis of non-Hermitian Kitaev chain
Our discussion so far has been focused on a simplified version of the non-Hermitian Kitaev chain, i.e., keeping the parameters $t_{1}$, $t_{2}$, $\mu$ real and setting $\varphi_{1}=\varphi_{2}=0$. We now study the non-Hermitian Kitaev chain by relaxing this constraint. With complex parameters, the analysis done so far becomes cumbersome. However, understanding the symmetries of the system makes the analysis easier. Here, following Ref. Kawabata_2022, we study the symmetries of the non-Hermitian Kitaev chain.
@@
In contrast to Hermitian systems, where the condition $H^{\dagger}=H$ (for Hamiltonian) ties together transpose and complex-conjugation operations, non-Hermitian Hamiltonians require that these operations be treated as independent when defining discrete symmetries. For example, when studying time-reversal symmetry (TRS) in a non-Hermitian system, one should address $H^{*}$ and $H^{T}$ separately, corresponding to TRS and its adjoint TRS. Similar statements hold for chiral-symmetry (CS) and particle-hole symmetry (PHS) Kawabata_2022. Considering arbitrary complex parameters, cf. Eq. (III) and by using Specht’s theorem Horn_Johnson_1985, we find the symmetry classification for the non-Hermitian Kitaev chain as shown in Tab. 1.
It is important to note that the particle-hole symmetry (PHS) always persists, with arbitrary complex parameters. The table (Tab. 1) presented here considers real $\mu$. For complex $\mu$, PHS still holds, and the only other possible symmetries are TRS and CS, see Appendix A. As a result of PHS, the Majorana zero modes exist and are robust across the entire parameter regime, Fig. 3a-c and Fig. 4a. However, the bulk modes exhibit NHSE with particle and hole localizing at opposite edges, Fig. 4b.
@@ III.4 Topological phase diagram for non-Hermitian Kitaev chain
We have seen that the topological phase transition point coincides for a finite open chain and an infinite or periodic chain in the non-Hermitian Kitaev chain. Additionally, we found that the zero-mode solution exists for a range of values of the chemical potential $\mu$, see Fig. 2. Hence, it is important to identify the topological and trivial phase in the non-Hermitian Kitaev chain, Eq. (III) as a function of the parameters. Therefore, we would like to have a topological invariant that can detect the phase, similar to the Hermitian version, cf. Sec. II.
Similar to the Hermitian Kitaev chain Alicea_2012, we begin by writing our Hamiltonian, Eq. (4) in the Pauli basis i.e., $H(k)=h_{\mathbb{I}}(k)\sigma_{\mathbb{I}}+h_{x}(k)\sigma_{x}+h_{y}(k)\sigma_{y}+h_{z}(k)\sigma_{z}$, where $\sigma_{\mathbb{I}}$ is the identity matrix and $\sigma_{x},\,\sigma_{y},\,\sigma_{z}$ are the Pauli matrices. The coefficients $h_{\mathbb{I}}(k),\,h_{x}(k),\,h_{y}(k)$ and $h_{z}(k)$ are given as follows,
@@
For $t_{1}+t_{2}\geq 2|\mu|$, we have $\nu=-1$, which is identified with the topological phase with two zero modes. In the case of $t_{1}+t_{2}<2|\mu|$, we have $\nu=+1$, which corresponds to the trivial phase, cf. Fig. 2a-c.

### v2

@@ I Introduction
Non-Hermitian quantum mechanics has engaged physicists for many years from various research areas, e.g. mathematical physics [6, 9, 5, 7, 51], complex space quantum mechanics [8, 30], and scattering theory [50]. More recently, researchers have been interested in understanding the quantum dynamics of open quantum systems, i.e. systems coupled with an environment [43, 19, 17, 24, 56, 49] for their potential application in quantum optics [55, 54] and quantum information theory [12, 38]. Moreover, non-Hermitian generalizations of lattice models provide an excellent testbed to study phenomena ubiquitous to non-Hermitian physics, such as complex spectra, the appearance of exceptional points and the non-Hermitian skin effect (NHSE) i.e., exponential localization of the eigenstates towards one edge (or boundary) [71, 60, 37, 29, 39, 74, 52, 42, 75, 31, 44, 68, 57, 27].
@@
The Hatano-Nelson model [25], a generalization of tight-binding Hamiltonian with non-reciprocal hopping amplitude is an example of a system that exhibits NHSE. An unusual feature of such systems is the high sensitivity of the spectrum to the boundary conditions. This manifests itself in the fact that under periodic boundary conditions, the spectrum is complex and the system is unstable, but surprisingly in a finite open chain, the spectrum is purely real. The eigenstates, instead of being Bloch waves, localize exponentially toward one edge, giving rise to the NHSE.
@@
In a set of seminal work [71, 61], it was shown that NHSE has a deep connection with topological properties of the system. It was shown that with non-reciprocal hoppings in the Su-Schrieffer-Heeger (SSH) model, the topological phase transition point in a finite open chain does not coincide with that of an infinite (or periodic) chain. The true transition point is governed by non-Bloch bulk-boundary correspondence, which is central to topological non-Hermitian systems. This discovery has inspired a slew of research at the intersection of non-Hermitian physics and topology [60, 20, 49, 75, 74, 52, 53, 10, 36, 72, 42, 67, 11, 31, 60, 37, 70, 45, 14, 77, 29, 39, 40, 66, 23, 44, 33, 32, 46, 3, 34, 16, 13, 47, 57, 48, 28].
@@
This naturally raises the question of whether such behavior is universal in one-dimensional systems. By considering a non-Hermitian Kitaev chain [35, 2, 73, 41, 69, 58, 59, 15, 4] with non-reciprocal nearest neighbor hopping amplitudes ($t_{1}$ and $t_{2}$) and asymmetric superconducting pairing terms ($\Delta_{1}e^{-i\varphi_{1}}$ and $\Delta_{2}e^{i\varphi_{2}}$), see Fig. 1, we show that this is not generally the case. For a non-Hermitian Kitaev chain, cf. Fig. 1, the topological phase transition points for finite open chain coincides with that of an infinite (or periodic) chain. By performing a symmetry analysis, we show that particle-hole symmetry (PHS) always exists across the entire parameter regime and thereby protects the Majorana modes from NHSE. However, we show that particle and hole bulk modes exhibit an equal and opposite skin effect, which is also a consequence of PHS. Finally, we present the phase diagram for the non-Hermitian Kitaev chain in the complex chemical potential ($\mu$) plane.
@@
This article is organized as follows. In section II, we briefly review the topological phase transition in the Kitaev chain. In section III, we introduce non-reciprocal hopping amplitudes and asymmetric superconducting pairing terms to construct a non-Hermitian Kitaev chain. Next, in section III.1, we numerically obtain the band structure, ground state and first excited state for a finite chain, and compare with infinite chain. In section III.2, we justify the robustness of Majorana zero modes to non-Hermiticity by considering the generalized Brillouin zone. In section III.3, we study the symmetry structure of the non-Hermitian Kitaev chain in the complex parameter space. Finally, in section III.4, we present topological phase diagrams with different parameter choices.
@@ II Kitaev chain
The one-dimensional (1D) Kitaev chain exhibits topological phase transition with two distinct phases in the parameter space. It is a tight-binding chain with $p$-wave superconducting pairing terms and hosts Majorana edge states in the topological phase [35]. We begin with a brief review of the 1D (Hermitian) Kitaev chain [2] before investigating its non-Hermitian extension. The Hamiltonian with $L$ lattice points, written in terms of particle creation ($c_{j}^{\dagger}$) and annihilation ($c_{j}$) operators on the $j^{\text{th}}$ site is given as follows,
@@
where $t$ is the hopping amplitude, $\Delta$ is the superconducting gap, $\varphi$ is the superconducting phase and $\mu$ is the chemical potential. This system exhibits a topological phase transition [65] as we tune the parameters ($t$ or $\mu$) of the Hamiltonian. The topological phase is marked by the existence of Majorana zero edge modes in a finite chain [35, 2, 1].
@@ III.1 Finite non-Hermitian Kitaev chain
where we have again assumed for simplicity $\mu,t_{1},t_{2}\in\mathbb{R}$ and $\varphi_{1}=\varphi_{2}=0$. Diagonalizing the Hamiltonian in Eq. (6) reveals a complex band structure, Fig. 2(b)-(c) and the presence of zero-energy modes in certain parameter regime, Fig. 2(a).
@@
The band structure shows that the zero mode solutions only appear for certain values of the chemical potential $\mu$, see Fig. 2(a), to be precise, the zero modes exist whenever we have $|\mu|<\frac{1}{2}(t_{1}+t_{2})$. This implies that the topological transition point derived using the spectrum of infinite chain, Eq. (III) coincides with the transition point obtained from the band structure of finite chain with OBC, cf. Fig. 2(a)-(b). In order to gain further insight, we look at the wave functions for the finite chain with OBC. With non-reciprocal hoppings ($t_{1}\neq t_{2}$), one would naively expect to see the non-Hermitian skin effect (NHSE) [71]. However, we find that the zero modes do not show NHSE and remain robust to non-reciprocity, Fig. 3(a)-(c). The NHSE only appears when we consider wavefunctions corresponding to excited states, Fig. 3(d). In section III.2 and section III.4, we present an explanation for the robustness of the zero modes to NHSE, Fig. 3(a)-(c), the presence of skin effect in excited states, Fig. 3(d).
@@ III.2 Generalized Brillouin zone
The NHSE is an exponential localization of wave functions and is the result of a modified Brillouin zone (BZ) in non-Hermitian systems [71]. Simplest way to obtain the modified Brillouin is to make the substitution $e^{ika}\to\beta$ (note that $a$ is the lattice constant and is set to one), where $|\beta|$ is determined by the non-reciprocities (or the non-Hermiticity) in the hopping parameters [71, 61]. Hamiltonian $H^{\text{NH}}$, Eq. (4) with the substitution $e^{ik}\to\beta$ becomes,
@@
From Eq. (9) we find that if we have $\beta_{0}$ as a solution, then $\beta_{0}^{-1}$ is also a solution. Hence, in general, we have four solutions to Eq. (9): $\beta_{1},\,\beta_{1}^{-1},\,\beta_{2},\,\beta_{2}^{-1}$. Additionally, one can show that the matrix, Eq. (8) can be null only at the boundary of the topological and trivial phase, hence the rank of that matrix is 1 at all points in the topological phase. Hence, corresponding to each $\beta$, there is only one zero mode eigenvector in the topological phase. Thus, the four roots of Eq. (9) will have four zero mode eigenvectors in the topological phase. Specifically, at $\mu=0$, we have $\beta_{2}=-\beta_{1}$. On physical grounds, we expect that the zero mode eigenvectors corresponding to $\beta_{1}(\beta_{1}^{-1})$ will be the same as $\beta_{1}(-\beta_{1}^{-1})$. Consequently, for $E=0$, the two solutions have inverse localization factors $\beta_{1}$ and $\beta_{1}^{-1}$, which means that the solution $\beta_{1}$ ($\beta_{1}^{-1}$) localizes to the right (left) [41, 69, 59]. The total probability density of the zero-energy states is given by $|\psi_{0}^{(1)}|^{2}+|\psi_{0}^{(2)}|^{2}$, where $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ are the normalized zero modes corresponding to factors $\beta_{1}$ and $\beta_{1}^{-1}$ respectively. This reciprocal pairing eliminates the skin effect for zero-energy states.
The existence of the reciprocal solution pair can also be understood by the symmetries of the Kitaev chain. Particle-hole symmetry enforces solutions to be reciprocal pairs [41]. In section III.3, we demonstrate that the particle-hole symmetry exists for arbitrary complex parameters in the Hamiltonian, thus protecting the Majorana edge modes, whenever they exist, for those parameter values. However, for any excited state (bulk mode), the solutions to $\text{det}(H^{\text{NH}}(\beta)-E)=0$ are not in a reciprocal pair. In fact, if $\beta_{1}$ is a solution corresponding to energy $E$, then $\beta_{1}^{-1}$ is a solution corresponding to energy E-E. Thus we expect an equal and opposite non-Hermitian skin effect for the particle and hole wavefunctions, cf. Fig. 4(b).
@@ III.3 Symmetry analysis of non-Hermitian Kitaev chain
Our discussion so far has been focused on a simplified version of the non-Hermitian Kitaev chain, i.e., keeping the parameters $t_{1}$, $t_{2}$, $\mu$ real and setting $\varphi_{1}=\varphi_{2}=0$. We now study the non-Hermitian Kitaev chain by relaxing this constraint. With complex parameters, the analysis done so far becomes cumbersome. However, understanding the symmetries of the system makes the analysis easier. Here, following Ref. [33], we study the symmetries of the non-Hermitian Kitaev chain.
@@
In contrast to Hermitian systems, where the condition $H^{\dagger}=H$ (for Hamiltonian) ties together transpose and complex-conjugation operations, non-Hermitian Hamiltonians require that these operations be treated as independent when defining discrete symmetries. For example, when studying time-reversal symmetry (TRS) in a non-Hermitian system, one should address $H^{*}$ and $H^{T}$ separately, corresponding to TRS and its adjoint TRS. Similar statements hold for chiral-symmetry (CS) and particle-hole symmetry (PHS) [33]. Considering arbitrary complex parameters, cf. Eq. (III) and by using Specht’s theorem [26], we find the symmetry classification for the non-Hermitian Kitaev chain as shown in Table 1.
It is important to note that the particle-hole symmetry (PHS) always persists, with arbitrary complex parameters. The Table 1 presented here considers real $\mu$. For complex $\mu$, PHS still holds, and the only other possible symmetries are TRS and CS, see Appendix A. As a result of PHS, the Majorana zero modes exist and are robust across the entire parameter regime, Fig. 3(a)-(c) and Fig. 4(a). However, the bulk modes exhibit NHSE with particle and hole localizing at opposite edges, Fig. 4(b).
@@ III.4 Topological phase diagram for non-Hermitian Kitaev chain
We have seen that the topological phase transition point coincides for a finite open chain and an infinite or periodic chain in the non-Hermitian Kitaev chain. Additionally, we found that the zero-mode solution exists for a range of values of the chemical potential $\mu$, see Fig. 2. Hence, it is important to identify the topological and trivial phase in the non-Hermitian Kitaev chain, Eq. (III) as a function of the parameters. Therefore, we would like to have a topological invariant that can detect the phase, similar to the Hermitian version, cf. section II.
Similar to the Hermitian Kitaev chain [2], we begin by writing our Hamiltonian, Eq. (4) in the Pauli basis i.e., $H(k)=h_{\mathbb{I}}(k)\sigma_{\mathbb{I}}+h_{x}(k)\sigma_{x}+h_{y}(k)\sigma_{y}+h_{z}(k)\sigma_{z}$, where $\sigma_{\mathbb{I}}$ is the identity matrix and $\sigma_{x},\,\sigma_{y},\,\sigma_{z}$ are the Pauli matrices. The coefficients $h_{\mathbb{I}}(k),\,h_{x}(k),\,h_{y}(k)$ and $h_{z}(k)$ are given as follows,
@@
For $t_{1}+t_{2}\geq 2|\mu|$, we have $\nu=-1$, which is identified with the topological phase with two zero modes. In the case of $t_{1}+t_{2}<2|\mu|$, we have $\nu=+1$, which corresponds to the trivial phase, cf. Fig. 2(a)-(c).